### v2

@@ 6.1 Testing for the CRS in a production function
Results with different basis functions are in Table 1. In all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
@@
We end this application with some remarks: (1) In addition to inference on CRS, our analysis also reveals other similarities with prior works. Table 2 shows that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model is computed by $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$. The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. Figures 1 (a)-(b) illustrates this with the semiparametric estimates with spatial weights $W_{n}^{d}$. They show that the output elasticity of capital is increasing in management expenses, while it is decreasing for labor. The decreasing elasticity of labor is a known empirical fact in this industry due to “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. (2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure 1 (c). Further discussions on these findings can be found in Li et al. (2002). (3) Our results are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion.
@@ 6.2 Testing the impact of the distance on house prices
Therefore, we estimate:
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
Results are listed in Table 3. All tests reject the null $H_{0}:\alpha=0$. In addition, testing with $\mathbb{W}_{4}$ leads to a rejection of $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. We further estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. As shown in Panel E of Table 3, the parametric method underestimates the spatial coefficient. Thus our results based on more general models complement those in Sun et al. (2014).
7.  Monte Carlo
@@ 7.1 Basic setting
To construct the SHAC statistics $\mathbb{W}_{2},\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we proceed as follows. Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 9 (a). [^4] From the symmetric circulant matrix $W_{d_{\lambda}}$, we build a graph and compute its pairwise shortest‐path distances in MATLAB, then apply cmdscale to generate $n\times 2$ coordinate matrix. We next compute the full $n\times n$ Euclidean distance matrix $D$ with the typical elements $d_{ij}$. For each m=1,,Mm=1,...,M, we add the actual distance matrix, $D_{m}^{*}$, used in practice with unobserved symmetric measurement noise. Its typical entries are $d_{ij,m}^{*}=d_{ij}+\nu_{ij,m},$ where $\nu_{ij,m}=1/2(\mu_{ij,m}+\mu_{ji,m})$ with $\mu_{ij,m}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. Define for each unit $i$ the $\ell$-th nearest‐neighbor distance by $d_{(i,\ell),m}=$ the $\ell$-th smallest entry of row $i$ of $D^{*}_{m}$, excluding the diagonal. We then set $d_{m}=\max_{1\leq i\leq n}d_{(i,\ell),m},$ so that every unit has at least $\ell$ neighbors within distance $d_{m}$. Collecting these over $m=1,\dots,M$ yields the $M\times 1$ vector $d_{m}$. We set $\delta=0$ for the null hypothesis and $\delta(\mathrm{x})=1-\mathrm{x}^{2}$ for any $\mathrm{x}\in R$ for the alternatives. We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the supplement. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[X,W_{1}X,...,W_{d_{\lambda}}X,\Psi]$. Furthermore, asy-p represents the standard-normal p-values, whereas chi-p represents the chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$, which is to improve testing performance in small samples. This may also help in empirical applications, although in both applications in the previous section critical values from either distribution yield the same results. We discuss size performance in this section and present power performance in the supplement, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
@@ 8.  Conclusion
Appendix
@@ Appendix A General Central Limit Theorem
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7 and $\mathscr{S}$ is a symmetric matrix. (b) The entries $s_{ij}$ of $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$. (c) $\mathscr{S}$ satisfies 1.
@@
Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 10. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
@@
Proof.
${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
@@
$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right$.
$\times\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
@@
$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right$.
$\times\left.\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
@@ Appendix B Proofs of Theorems
Proof of Theorem 2.1.
Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}\\ P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
@@
where $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$ with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, and $A_{1}=r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$ and $A_{2}=2r_{\delta}^{\prime}\mathcal{M}_{1}\epsilon$. Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 1 and 4. A similar argument holds for $A_{2}$ because it has (conditionally) zero mean and variance $4r_{\delta}^{\prime}\mathcal{M}_{1}E(\epsilon\epsilon^{\prime})\mathcal{M}_{1}r_{\delta}=4r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$, by idempotence of $\mathcal{M}_{1}$. Thus $\frac{A_{1}}{\sqrt{d_{\alpha}}}$ and $\frac{A_{2}}{\sqrt{d_{\alpha}}}$ are both $o_{p}(1)$ and establishing the theorem. ∎
Proof of Theorem $2.$
This is a special case of Theorem 3.2 and therefore omitted. ∎
Proof of Lemma 3.1.
This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 3.1.
$\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
@@
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, $\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and $\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 3.2.
The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1. Checking Assumption A follows exactly as in the proof of Theorem 5.2 below and is omitted. ∎
Proof of Lemma 4.1.
This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 4.1.
Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$. Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big\|r_{G\delta}\big\|^{2}=o_{p}(1)$ by Assumption 12. Under $H_{0}:\alpha=0$,

### v3

@@ 6.1 Testing for the CRS in a production function
For the test statistics, $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{d_{\mu}})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for k=1,,hk=1,...,h, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function k=1,,h/2k=1,...,h/2 for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function. For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and $l=1,...,d_{\tau}$, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$. Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$. For $\mathbb{W}_{4}$, we use $\phi_{im}(z)=\frac{1}{d_{\mu}}[\frac{2}{\pi}\tanh(z_{i})]^{m}$ as the polynomial function and $\phi_{im}(z)=\frac{1}{d_{\mu}}\sin(\frac{z_{i}}{2m})$ as the trigonometric function for $m=1,...,d_{\mu}$. Then, $\mathfrak{H}_{im}=(\sum_{j=1}^{n}w_{ij}y_{j})\phi_{im}(z)$, where $w_{ij}$ denotes the $(i,j)$-th entry of either $W_{n}^{p}$ or $W_{n}^{d}$. For the IVs, set $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X]$ when the specification uses $W_{n}^{p}$, and $K_{1}=K_{3}=[X,\Psi,W_{n}^{d}X]$ when it uses $W_{n}^{d}$, for $\mathbb{W}_{1}$, $\mathbb{W}_{2}$, and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{d_{\tau}}X]$ with respect to $\mathbb{W}_{3}$. We set $d_{\mu}=d_{\tau}=h$, and the results with different basis functions are in Table 2. In all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). These non-rejection results are economically meaningful. One can conclude that the CRS technology is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
@@
We end this application with some remarks: (1) In addition to inference on CRS, our analysis also reveals other similarities with prior works. Table 1 shows that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model are computed by summing $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$.The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. Figures S1 (a)-(b) in the online appendix illustrate this with the semiparametric estimates with spatial weights $W_{n}^{d}$. They show that the output elasticity of capital is increasing in management expenses, while it is decreasing for labor. The decreasing elasticity of labor is a known empirical fact in this industry due to “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. (2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure S1 (c) in the online appendix. Further discussions on these findings can be found in Li et al. (2002). (3) Our results are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion.
Without $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
$\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
$\hat{\delta}_{2}$ 0.3664 0.3212 0.3691 0.2930 0.3502 0.3220
$\hat{\delta}_{2}$ 0.3517 0.3099 0.3381 0.3040 0.3377 0.2816
$\hat{\delta}_{2}$ 0.3089 0.3395 0.4345 0.3252 0.3471 0.3025
$\hat{\delta}_{2}$ 0.3121 0.2702 0.3002 0.3043 0.3649 0.3199
@@ 6.2 Testing the impact of the distance on house prices
Let $d_{\tau}=d_{\mu}=h$, we estimate:
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}\mathfrak{H}_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Results are listed in Table 3. All tests reject the null $H_{0}:\alpha=0$. In addition, testing with $\mathbb{W}_{4}$ leads to a rejection of $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. We further estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. As shown in Panel E of Table 3, the parametric method underestimates the spatial coefficient. Thus, our results based on more general models complement those in Sun et al. (2014).
7 Monte Carlo Simulation
@@ 7.1 Basic setting
To construct the SHAC statistics $\mathbb{W}_{2},\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we proceed as follows. Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 3.4 (a). [^4] From the symmetric circulant matrix $W_{d_{\lambda}}$, we build a graph and compute its pairwise shortest‐path distances in MATLAB, then apply cmdscale to generate $n\times 2$ coordinate matrix. We next compute the full $n\times n$ Euclidean distance matrix $D$ with the typical elements $d_{ij}$. For each m=1,,Mm=1,...,M, we add the actual distance matrix, $D_{m}^{*}$, used in practice with unobserved symmetric measurement noise. Its typical entries are $d_{ij,m}^{*}=d_{ij}+\nu_{ij,m},$ where $\nu_{ij,m}=1/2(\mu_{ij,m}+\mu_{ji,m})$ with $\mu_{ij,m}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. Define for each unit $i$ the $\ell$-th nearest‐neighbor distance by $d_{(i,\ell),m}=$ the $\ell$-th smallest entry of row $i$ of $D^{*}_{m}$, excluding the diagonal. We then set $d_{m}=\max_{1\leq i\leq n}d_{(i,\ell),m},$ so that every unit has at least $\ell$ neighbors within distance $d_{m}$. Collecting these over $m=1,\dots,M$ yields the $M\times 1$ vector $d_{m}$. We set $\delta=0$ for the null hypothesis and $\delta(\mathrm{x})=1-\mathrm{x}^{2}$ for any $\mathrm{x}\in R$ for the alternatives. We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the online appendix. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[X,W_{1}X,...,W_{d_{\lambda}}X,\Psi]$. Furthermore, asy-p represents the standard-normal p-values, whereas chi-p represents the chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$, which is to improve testing performance in small samples. This may also help in empirical applications, although in both applications in the previous section critical values from either distribution yield the same results. We discuss size performance in this section and present power performance in the online appendix, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
@@ 8 Conclusion
A Supplement to
Wald inference on varying coefficients
Abhimanyu Gupta
Department of Economics, Queen's University, Dunning Hall, 94 University Avenue, Kingston, Ontario K7L 3N6, Canada and Department of Economics, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ, UK. Email: abhimanyu.g@queensu.ca.
Xi Qu
Department of Economics, Antai College of Economics and Management, Shanghai Jiao Tong University, 1954 Huashan Road, Shanghai, 200030, China PRC. Email: xiqu@sjtu.edu.cn.
Sorawoot Srisuma
Department of Economics, National University of Singapore, 1 Arts Link, 117570, Singapore.
Jiajun Zhang
International Business School, Shanghai University of International Business and Economics, 201620, China PRC. E-mail: jiajun30@suibe.edu.cn.
@@ Appendix A General Central Limit Theorem
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 3.2 and $\mathscr{S}$ is a symmetric matrix. (b) The entries $s_{ij}$ of $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$. (c) $\mathscr{S}$ satisfies 1.
@@
Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 3.5. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
@@
(Proof). ${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
@@
$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\times\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
@@
$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\times\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
@@ Appendix B Proofs of Theorems
(Proof).[Proof of Theorem 2.1] Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
@@
where $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$ with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, and $A_{1}=r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$ and $A_{2}=2r_{\delta}^{\prime}\mathcal{M}_{1}\epsilon$. Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 2.1 and 2.4. A similar argument holds for $A_{2}$ because it has (conditionally) zero mean and variance $4r_{\delta}^{\prime}\mathcal{M}_{1}E(\epsilon\epsilon^{\prime})\mathcal{M}_{1}r_{\delta}=4r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$, by idempotence of $\mathcal{M}_{1}$. Thus $\frac{A_{1}}{\sqrt{d_{\alpha}}}$ and $\frac{A_{2}}{\sqrt{d_{\alpha}}}$ are both $o_{p}(1)$ and establishing the theorem. $\blacksquare$
(Proof).[Proof of Theorem 2.2] This is a special case of Theorem 3.2 and therefore omitted. $\blacksquare$
(Proof).[Proof of Lemma 3.1] This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. $\blacksquare$
(Proof).[Proof of Theorem 3.1] $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
@@
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, $\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and $\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. $\blacksquare$
(Proof).[Proof of Theorem 3.2] The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1. Checking Assumption A follows exactly as in the proof of Theorem 5.2 below and is omitted. $\blacksquare$
(Proof).[Proof of Lemma 4.1] This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. $\blacksquare$
(Proof).[Proof of Theorem 4.1] Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$. Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big\|r_{G\delta}\big\|^{2}=o_{p}(1)$ by Assumption 4.2. Under $H_{0}:\alpha=0$,